Fill out the table and determine the value of the following quantities, Using the following p.m.f.

f(x)= (x+1)2/91, x=0,1,2,3,4,5

i.Probability Table:

x

f(x)

xf(x)

0

1

2

3

4

5

ii . What is the E[X]?

iii. What is the E[x2]?

iv.  What is Var(x)?
Answer #1

SOLUTION:

From given data,

Fill out the table and determine the value of the following quantities, Using the following p.m.f.

f(x)= (x+1)2/91, x=0,1,2,3,4,5

i.Probability Table:

x f(x) x*f(x) x2*f(x)
0 0.0109 0 0
1 0.0439 0.0439 0.0439
2 0.0989 0.1978 0.3956
3 0.1758 0.5274 1.5822
4 0.2747 1.0988 4.3952
5 0.3956 1.978 9.89

ii . What is the E[X]?

E[X] = x*f(x) = 0+0.0439+0.1978+0.5274+1.0988+1.978 = 3.8459

E[X] = 3.8459

iii. What is the E[x2]?

E[X2] = x2*f(x) = 0+0.0439+0.3956+1.5822+4.3952+9.89 = 16.3069

E[X2] = 16.3069

iv.  What is Var(x)?

Var(x) = E[X2] - (E[X])2

Var(x) = 16.3069 - (3.8459)2

Var(x) = 16.3069 - 14.79094681

Var(x) = 1.5159

Please thumbs-up / vote up this answer if it was helpful. In case of any problem, please comment below. I will surely help. Down-votes are permanent and not notified to us, so we can't help in that case.